OpenStudy (toxicsugar22):

acetic acid pKa=4.76, ka__________ M

OpenStudy (asevilla5):

http://www.webqc.org/phsolver.php this might help you not sure

OpenStudy (kainui):

\[\LARGE pK_a=-\log_{10}(K_a)\]

OpenStudy (kkutie7):

\[10^{-pKa}\]

OpenStudy (kkutie7):

someone gave you the equation as \[pKa=-log(K_{a})\] solve for Ka \[-pKa=log(K_{a})\] \[10^{-pKa}=K_{a}\] you know pKa plug it in and solve

OpenStudy (toxicsugar22):

is it 1.737800829 10^-5

OpenStudy (kkutie7):

yes

OpenStudy (kkutie7):

I'm not sure sorry

OpenStudy (australopithecus):

pKa = -Log(Ka)

OpenStudy (australopithecus):

toxicsugar22 if you dont trust her alegbra you are can always look up rules for manipulating logarithms

OpenStudy (australopithecus):

I will tell you though that it is correct but you should know its correct.

OpenStudy (australopithecus):

your algebra is wrong

OpenStudy (australopithecus):

Look at kkutie7s answer she does the algebra correctly, http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (kkutie7):

you get 10^(-pKa) because you need to get rid of the log on the right hand side so you do 10^ to each side to just get Ka. as for sig figs... I can't remember the rules right now, but I usually put it as 1.74*10^(-5)

OpenStudy (toxicsugar22):

is anyone formalize with the Henderson hasselbach formula

OpenStudy (australopithecus):

Yes it is very easy

OpenStudy (toxicsugar22):

cna you apply this to thate qation and find ka from that

OpenStudy (toxicsugar22):

ph=pka+log{a-/{HA}

OpenStudy (toxicsugar22):

do i get the same anwer and how would i do that

OpenStudy (toxicsugar22):

or i cant use this fromula for acetic acid pKa=4.76, ka__________ M

OpenStudy (australopithecus):

I gave you the formula pKa = -Log(Ka) Ka = 10^(-pKa)

OpenStudy (toxicsugar22):

1.737800829 10^-5

OpenStudy (australopithecus):

http://www.wolframalpha.com/input/?i=Ka+%3D+10%5E%28-4.76%29

OpenStudy (australopithecus):

that is your answer